# Improved Global Landscape Guarantees for Low-rank Factorization in Synchronization

### v1

@@ 2 Main results
which follows from dropping all the negative terms in (2.2) and using $u\leq w\leq v\leq 1.$ This upper bound is clean because it does not involve $\alpha.$ Combining this upper bound of $G(x,\tau)$ with the master theorem and Corollary 2.2, we have the following result.
@@
Note that the higher the noise level, the larger the condition number of $\bm{L}$, and thus a larger $\alpha_{G}(p,d)$ means the optimization landscape is more robust against the noise in the data for any given pair of $(p,d)$. Apparently, this has already improved the existing work by the author [26] because for $\tau=1$, then
@@ 2.1 Exact computation of $\alpha_{G}(p,d)$
$G(x,\tau)=\begin{dcases}2q(\tau)+2d\tau^{2}-\dfrac{1}{xd},&xdq(\tau)\geq\max\left\{1,\frac{1}{h(\tau)}\right\},\\ xdq^{2}(\tau)+2d\tau^{2},&2-h(\tau)\leq xdq(\tau)\leq 1,\\ 2q(\tau)+2d\tau^{2}-q(\tau)h(\tau),&{\rm otherwise}.\end{dcases}$ (2.7)
@@
For $d=1$, then
[…]
@@
[…]
@@
[…]
@@
where $r(\tau)=p-1.$ Then
[…]
The maximum of $\alpha_{G}(p,1,\tau)$ over $\tau$ is attained at $\tau=0$ and $\alpha_{G}(p,1)=p.$
For $d=1$, Theorem 2.6 reproduces the remarkable result in [18], and also [30]. For $d\geq 2$, $\alpha_{G}(p,d)$ does not have an explicit form. We briefly discuss how this search for $\alpha_{G}(p,d,\tau)$ works. We take (2.11) as an example: for any $\tau$ with $h(\tau)\leq 1$, we first solve a quadratic equation for the largest $\alpha_{*}$ satisfying:
$\frac{r(\tau)}{\alpha}\geq 2(q(\tau)+d\tau^{2})-\frac{1}{(1-\alpha^{-1})d}.$
Then we check if $\alpha_{*}$ satisfies the constraints $1-\alpha_{*}^{-1}\geq(dq(\tau)h(\tau))^{-1}$. If it satisfies, then we set $\alpha_{G}(p,d,\tau)=\alpha_{*}$; otherwise, we go to the next case. We later show that $\alpha_{G}(p,d)$ is guaranteed to attain in the first case of (2.10) and (2.11) in Theorem 2.8 if $p$ is slightly larger than $d$. This procedure does not yield a simple closed form but we can easily compute it numerically, as shown in the table.
@@
Now we provide the examples of $\alpha_{M}(p,d,\tau)$ for $\tau=0$ and 1. In particular, the following corollary implies the author’s earlier work [26] for $\tau=1$.
@@
where the maximizer is the larger root of the quadratic equation
@@
In other words, under (2.12), $\alpha_{G}(p,d)$ is attained in the first case of either (2.10) or (2.11).
@@ 2.2 Tightness of the bound and future work
is a family of diagonal matrices of size $d\times d$ with $\{\pm 1\}$ diagonal entries and
@@ 2.4 Roadmap of the proof
The idea of constructing a feasible dual variable in (2.16) is to first simplify some constraints. Unlike the earlier works [26, 18] in which a specific choice of $\dot{\bm{S}}_{i}$ is chosen, we will consider a family of tangent vectors and try to find the optimal one for any given pair of $(p,d).$ We first introduce this family of tangent vectors, and present a few important quantities that will be used in the construction of feasible dual variables.
@@
The properties of ${\cal L}_{\bm{S}_{i}}(\cdot)$, ${\cal L}_{\bm{S}}(\cdot)$, ${\cal P}_{\bm{S}_{i}}(\cdot)$, and ${\cal P}_{\bm{S}}(\cdot)$ are discussed in Section 3.2. In particular, we define
@@
$\langle\bm{L},\bm{X}_{\perp}\rangle-\frac{\alpha\langle\bm{L},\bm{S}\bm{S}^{\top}\rangle}{\langle\bm{P}_{\perp},\bm{S}\bm{S}^{\top}\rangle}<0$
@@ 3.1 Formulation of the condition number minimization problem
Therefore, we arrive at the optimization problem 2.14. ∎
@@ 3.4 Reduction of the dual program $\alpha$ and its solution
where $\alpha_{M}^{\prime}(p,d,\tau)$ is the derivative w.r.t. $\tau.$ It is easy to check that the maximum is attained at the larger root to $\alpha_{M}^{\prime}(p,d,\tau)=0$, i.e.,
@@ 3.5 A refined estimation of $\alpha$
Thus $\alpha$ is the larger root to the quadratic equation that satisfies $(1-\alpha^{-1})dq(\tau)>1.$ If such a root does not satisfy $(1-\alpha^{-1})dq(\tau)>1$, it means a feasible $\alpha$ does not exist in this region.
@@
which corresponds to the second case of (2.10). With this choice of $\tau_{\epsilon}$, $G(x,\tau_{\epsilon})=xdq^{2}(\tau_{\epsilon})+2d\tau_{\epsilon}^{2}$ and
@@
Next, we will show that $\max_{0\leq\tau\leq 1}\alpha_{\widetilde{M}}(p,d,\tau)$ satisfies the constraint in the first case of (2.10) or (2.11). We first find the maximizer of the estimation $\alpha_{\widetilde{M}}(p,d,\tau)$ which provides the best lower bound:
@@
Let $\tau_{*}$ be the positive root to the quadratic equation. Then
@@
is a lower bound of $\sup_{0\leq\tau\leq 1}\alpha_{G}(p,d,\tau)$ provided that the first constraint of (2.10) or (2.11) holds, i.e.,

### v2

@@ 2 Main results
which follows from dropping all the negative terms in (2.2) and using $u\leq w\leq v\leq 1.$ This upper bound is clean because it does not involve $x.$ Combining this upper bound of $G(x,\tau)$ with the master theorem and Corollary 2.2, we have the following result.
@@
Note that the higher the noise level, the larger the condition number of $\bm{L}$, and thus a larger $\alpha_{G}(p,d)$ means the optimization landscape is more robust against the noise in the data for any given pair of $(p,d)$. Apparently, this has already improved the existing work [26] by the author because for $\tau=1$, then
@@ 2.1 Exact computation of $\alpha_{G}(p,d)$
$G(x,\tau)=\begin{dcases}2q(\tau)+2d\tau^{2}-\dfrac{1}{xd},&{\rm if}~~xdq(\tau)\geq\max\left\{1,\frac{1}{h(\tau)}\right\},\\ xdq^{2}(\tau)+2d\tau^{2},&{\rm if}~~2-h(\tau)\leq xdq(\tau)\leq 1,\\ 2q(\tau)+2d\tau^{2}-q(\tau)h(\tau),&{\rm otherwise}.\end{dcases}$ (2.7)
@@
For $d=1$, then $\alpha_{G}(p,1,\tau)$ is given as follows.
  1. 1.
    $\alpha_{G}(p,1,\tau)=\alpha$ if $\alpha=\frac{p-1+(1-\tau^{2})^{2}}{2\tau^{2}+(1-\tau^{2})^{2}}$ satisfies
    $(1-\alpha^{-1})(1-\tau^{2})^{2}+2\tau^{2}\geq\frac{p-1}{p};$ (2.9)
  2. 2.
    otherwise, $\alpha_{G}(p,1,\tau)=p$ if $\alpha=p$ satisfies
    $(1-\alpha^{-1})(1-\tau^{2})^{2}+2\tau^{2}\geq\frac{p-1}{p}.$ (2.10)
@@
  1. 1.
    $\alpha_{G}(p,d,\tau)=\alpha$ if $\alpha$ is the larger root of the equation $\frac{r(\tau)}{\alpha}=2(q(\tau)+d\tau^{2})-\frac{1}{(1-\alpha^{-1})d}$, and satisfies
    $1-\alpha^{-1}\geq\frac{1}{dq(\tau)};$ (2.11)
  2. 2.
    otherwise, $\alpha_{G}(p,d,\tau)=\alpha$ if $\alpha=\frac{dq^{2}(\tau)+r(\tau)}{dq^{2}(\tau)+2d\tau^{2}}$ satisfies
    $\frac{2-h(\tau)}{dq(\tau)}\leq 1-\alpha^{-1}\leq\frac{1}{dq(\tau)};$ (2.12)
  3. 3.
    otherwise, $\alpha_{G}(p,d,\tau)=\alpha$ if $\alpha=\frac{r(\tau)}{2(q(\tau)+d\tau^{2})-q(\tau)h(\tau)}$ satisfies
    $1-\alpha^{-1}\leq\frac{2-h(\tau)}{dq(\tau)}.$ (2.13)
@@
  1. 1.
    $\alpha_{G}(p,d,\tau)=\alpha$ if $\alpha$ is the larger root of the equation $\frac{r(\tau)}{\alpha}=2(q(\tau)+d\tau^{2})-\frac{1}{(1-\alpha^{-1})d}$, and satisfies
    $1-\alpha^{-1}\geq\frac{1}{dq(\tau)h(\tau)};$ (2.14)
  2. 2.
    otherwise, $\alpha_{G}(p,d,\tau)=\alpha$ if $\alpha=\frac{r(\tau)}{2(q(\tau)+d\tau^{2})-q(\tau)h(\tau)}$ satisfies
    $1-\alpha^{-1}\leq\frac{1}{dq(\tau)h(\tau)}.$ (2.15)
@@
where $r(\tau)=p-1.$ Then we have $\alpha_{G}(p,1,\tau)$ in (2.9)-(2.10). The maximum of $\alpha_{G}(p,1,\tau)$ over $\tau$ is attained at $\tau=0$ and $\alpha_{G}(p,1)=p.$
For $d=1$, Theorem 2.6 reproduces the remarkable result in [18], and also [30]. For $d\geq 2$, $\alpha_{G}(p,d)$ does not have an explicit form. We briefly discuss how this search for $\alpha_{G}(p,d,\tau)$ works. We take (2.14)-(2.15) as an example: for any $\tau$ with $h(\tau)\leq 1$, we first solve a quadratic equation for the larger $\alpha_{*}$ satisfying:
$\frac{r(\tau)}{\alpha}=2(q(\tau)+d\tau^{2})-\frac{1}{(1-\alpha^{-1})d}.$
Then we check if $\alpha_{*}$ satisfies the constraints $1-\alpha_{*}^{-1}\geq(dq(\tau)h(\tau))^{-1}$. If it satisfies, then we set $\alpha_{G}(p,d,\tau)=\alpha_{*}$; otherwise, we go to the next case. We later show that $\alpha_{G}(p,d)$ is guaranteed to attain in the case (2.11) or (2.14) in Theorem 2.8 if $p$ is slightly larger than $d$. This procedure does not yield a simple closed form but we can easily compute it numerically, as shown in the table.
@@
Now we provide the examples of $\alpha_{G}(p,d,\tau)$ for $\tau=0$ and 1. In particular, the following corollary implies the author’s earlier work [26] for $\tau=1$.
@@
where the maximizer $\tau_{*}$ is the larger root of the quadratic equation
@@
In other words, under (2.16), $\alpha_{G}(p,d)$ is attained in the case (2.11) or (2.14).
@@ 2.2 Tightness of the bound and future work
is a family of diagonal matrices of size $d\times d$ with diagonal entries $\{\pm 1\}$ and
@@ 2.4 Roadmap of the proof
The idea of constructing a feasible dual variable in (2.20) is to first simplify some constraints. Unlike the earlier works [18, 26, 33] in which a specific choice of $\dot{\bm{S}}_{i}$ is chosen, we will consider a family of tangent vectors and try to find the optimal one for any given pair of $(p,d).$ We first introduce this family of tangent vectors, and present a few important quantities that will be used in the construction of feasible dual variables.
@@
The properties of ${\cal L}_{\bm{S}_{i},\tau}(\cdot)$, ${\cal L}_{\bm{S},\tau}(\cdot)$, ${\cal P}_{\bm{S}_{i},\tau}(\cdot)$, and ${\cal P}_{\bm{S},\tau}(\cdot)$ are discussed in Section 3.2. In particular, we define
@@
$\langle\bm{L},\bm{X}_{\perp}\rangle-\frac{\alpha\langle\bm{L},\bm{S}\bm{S}^{\top}\rangle}{\langle\bm{P}_{\perp},\bm{S}\bm{S}^{\top}\rangle}\leq 0$
@@ 3.1 Formulation of the condition number minimization problem
Therefore, we arrive at the optimization problem 2.18. ∎
@@ 3.4 Reduction of the dual program $\alpha$ and its solution
where $\alpha_{M}^{\prime}(p,d,\tau)$ is the derivative w.r.t. $\tau.$ It is easy to check that the maximum is attained at the larger root of $\alpha_{M}^{\prime}(p,d,\tau)=0$, i.e.,
@@ 3.5 A refined estimation of $\alpha$
Thus $\alpha$ is the larger root of the quadratic equation that satisfies $(1-\alpha^{-1})dq(\tau)>1.$ If such a root does not satisfy $(1-\alpha^{-1})dq(\tau)>1$, it means a feasible $\alpha$ does not exist in this region.
@@
which corresponds to (2.12). With this choice of $\tau_{\epsilon}$, $G(x,\tau_{\epsilon})=xdq^{2}(\tau_{\epsilon})+2d\tau_{\epsilon}^{2}$ and
@@
Next, we will show that $\max_{0\leq\tau\leq 1}\alpha_{\widetilde{M}}(p,d,\tau)$ satisfies the constraint in of (2.11) or (2.14). We first find the maximizer of the estimation $\alpha_{\widetilde{M}}(p,d,\tau)$ which provides the best lower bound:
@@
Let $\tau_{*}$ be the positive root of the quadratic equation. Then
@@
is a lower bound of $\sup_{0\leq\tau\leq 1}\alpha_{G}(p,d,\tau)$ provided that the constraint in (2.11) or (2.14) holds, i.e.,